A certain second-order reaction (B→products) has a rate constant of 1.55×10−3 M−1⋅s−1 at 27 ∘C and an initial half-life of 248 s . What is the concentration of the reactant B after one half-life?
Answer #1

1st find the initial concentration

[A]o = 1/(k*half life)

= 1/(1.55*10^-3 M-1.s-1 * 248 s)

= 2.601 M

After half life, concentration decreases to half the value it was before the hallf life

So, after half life,

concentration = initial concentration / 2

= (2.601) / 2

= 1.30 M

Answer: 1.30 M